Q1)Solve the equation: (Cosx)^7 +(sinx )^4=1.
I solved this question 2 times but each time got 2 different results and they both were wrong. But i I know that there were no mathematical errors in my solution then why i am getting a wrong answer?

Grade:12th pass

1 Answers

mkmjnj
5 years ago
let  a=cosx  and   b= sinx
we know    1-b2 = a2
a7 + b4 = 1
a7 = 1-b4
a7 = (1-b2 )(1+b2 )
a7 = a2 (1+b2)
take RHS term to LHS
a2 (a5 – 1--b2 )=0
therefore either
  • a2 =0  ;    cos2x = 0 ;    or
  • a5 = 1+b2 ;  cos5x = 1+ sin2x;    therefore     cosx =1    and        sinx = 0 ;    this is because RHS >= 1   but LHS
  • bro you can now easily solve the above trigonometric equations and get the desired answer.                      hope the given solution works